Ty is one-third as old as his father.Five years ago his father was four times as old as Ty.How old is each now?
Answer by ankor@dixie-net.com(22740)

Ty is one-third as old as his father.Five years ago his father was four times as old as Ty.How old is each now?
:
let f = fathers age noe
let t = Ty's age now
:
write an equation for each statement:
"Ty is one-third as old as his father."
f = 3t
:
"
f - 5 = 4(t-5)
f-5 = 4t - 20
f = 4t - 20 + 5
f = 4t - 15
:
How old is each now?
Substitute 3t for f in the above eq
3t = 4t - 15
3t - 4t = -15
-t = -15
t = 15 yrs is Ty's present age
then
3 * 15 = 45 yrs is father's present age
:
:
Check solution in the statement:
"Five years ago his father was four times as old as Ty."
45 - 5 = 4(15 - 5)